# Central Limit Theorems for Asynchronous Averaged Q-Learning

### v1

@@ Appendix C Supporting Lemmas
In this section we present several supporting lemmas. Lemma C.1 and C.5 analyze Term (1) and $\Psi_{i}^{K}$ appeared in the proof of Theorem 3.1. Next, by leveraging the results in [Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam], Lemma C.7 gives a non-asymptotic convergence rate for $\Delta_{k}=Q_{k}-Q^{*}$ under asynchronous updates. Lastly, Lemma C.8 provides a Lipschitz property for the operator $F(\cdot,\cdot)$ defined in eq. 2.
@@
Let $\alpha_{i}=\frac{\alpha}{(i+b)^{\beta}}$ for some problem-dependent constants $\alpha,b>0$ and $\beta\in(0,1)$. Then the following bounds hold:
@@
The analysis of polynomial step sizes has been well studied in prior work (see, e.g., [Polyak and Juditsky(1992), Srikant(2024), Li et al.(2023)Li, Yang, Liang, Zhang, and Jordan]). However, due to a slightly modified choice of the step-size and the different update rule in the asynchronous setting, we provide a complete proof for the sake of completeness. Recall that $\alpha_{i}=\alpha{(i+b)^{-\beta}}$ and $\rho:=\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\min}p(s,a)$. We now have
@@
Next, we prove the second part. Recall $\Psi_{i}^{K}=\alpha_{i}\sum_{k=i+1}^{K}\left(\prod_{j=i+1}^{k-1}(I-\alpha_{j}A)\right)$. Since $A^{-1}=\alpha_{i}^{-1}(I-(I-\alpha_{i}G))$, we have
@@
Let $\alpha_{i}=\frac{\alpha}{(i+b)^{\beta}}$ for some problem-dependent constants $\alpha,b>0$ and $\beta\in(0,1)$. It follows that
@@
The following lemmas provide finite-sample convergence guarantees of asynchronous Q-learning and Lipschitzness of the operator $F(\cdot,s)$ [Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam].
Lemma C.7 (Theorem B.1 in [Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam]).
Let $\alpha_{i}=\frac{\alpha}{(i+b)^{\beta}}$ for some problem-dependent constants $\alpha,b>0$ and $\beta\in(0,1)$. For the Q-learning updates in eq. 1, under Assumption 2, we have $\mathbb{E}\|Q_{k}-Q^{*}\|_{\infty}\leq O\left(\sqrt{\frac{t_{k}}{(1-\gamma)^{2}\rho^{2}k}}\right)$, where $t_{k}=O(\log(1/t))$ denotes the mixing time.
Lemma C.8 (Proposition 3.1 in [Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam]).

### v2

@@ Appendix C Supporting Lemmas
In this section we present several supporting lemmas. Lemma C.1 and C.5 analyze Term (1) and $\Psi_{i}^{K}$ appeared in the proof of Theorem 3.1. Next, by leveraging the results in Chen et al. (2021), Lemma C.7 gives a non-asymptotic convergence rate for $\Delta_{k}=Q_{k}-Q^{*}$ under asynchronous updates. Lastly, Lemma C.8 provides a Lipschitz property for the operator $F(\cdot,s)$ defined in eq. 2.
@@
Let $\alpha_{i}=\alpha{(i+b)^{-\beta}}$ for some problem-dependent constants $\alpha,b>0$ and $\beta\in(0,1)$. Then the following bounds hold:
@@
The analysis of polynomial step sizes has been well studied in prior work (see, e.g., Polyak and Juditsky (1992); Srikant (2024); Li et al. (2023)). However, due to a slightly modified choice of the step-size and the different update rule in the asynchronous setting, we provide a complete proof for the sake of completeness. Recall that $\alpha_{i}=\alpha{(i+b)^{-\beta}}$ and $\rho:=\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\min}p(s,a)$. We now have
@@
Next, we prove the second part. Recall $\Psi_{i}^{K}=\alpha_{i}\sum_{k=i+1}^{K}\left(\prod_{j=i+1}^{k-1}(I-\alpha_{j}A)\right)$. Since $A^{-1}=\alpha_{i}^{-1}(I-(I-\alpha_{i}A))$, we have
@@
Let $\alpha_{i}=\alpha{(i+b)^{-\beta}}$ for some problem-dependent constants $\alpha,b>0$ and $\beta\in(0,1)$. It follows that
@@
The following lemmas provide finite-sample convergence guarantees of asynchronous Q-learning and Lipschitzness of the operator $F(\cdot,s)$ (Chen et al., 2021).
Lemma C.7 (Theorem B.1 in Chen et al. (2021)).
Let $\alpha_{i}=\alpha{(i+b)^{-\beta}}$ for some problem-dependent constants $\alpha,b>0$ and $\beta\in(0,1)$. For the Q-learning updates in eq. 1, under Assumption 2, we have $\mathbb{E}\|Q_{k}-Q^{*}\|_{\infty}\leq O\left(\sqrt{\frac{t_{k}}{(1-\gamma)^{2}\rho^{2}k}}\right)$, where $t_{k}=O(\log(1/\alpha_{k}))$ denotes the mixing time.
Lemma C.8 (Proposition 3.1 in Chen et al. (2021)).